Filer Manufacturing has 4,372,396 shares of common stock outstanding. The current share price is $63.96, and the book value per share is $3.69. Filer Manufacturing also has two bond issues outstanding. The first bond issue has a face value of $55,719,562, has a 0.07 coupon, matures in 7 years and sells for 80 percent of par. The second issue has a face value of $48,080,635, has a 0.07 coupon, matures in 18 years, and sells for 97 percent of par. What is Filer's weight of equity on a market value basis? Enter the answer with 4 decimals (e.g. 0.2345)
Answer #1

Total equity = number of shares*price per share = 4,372,396*63.96 = 279,658,448.16

Market value of Bond 1 (MV1) = current price = 80% of par = 80%*55,719,562 = 44,575,649.60

Market value of Bond 2 (MV2) = current price = 97% of par = 97%*48,080,635 = 46,638,215.95

Total market capital (TMC) = total equity + MV1 + MV2 = 279,658,448.16 + 44,575,649.60 + 46,638,215.95 = 370,872,313.71

Weight of equity (on market value basis) = total equity/TMC = 279,658,448.16/370,872,313.71 = 0.7541 (or 75.41%)